OpenStudy (anonymous):

Pedro has created the function f(x) = (4x-3)/2 to represent the number of assignments he has completed, where x represents the number of weeks in the course. Pedro discovers that, using the inverse function to solve for x = 30, he can predict when he will have 30 assignments completed. Explain to Pedro how to accomplish this, using complete sentences.

OpenStudy (freckles):

y=(4x-3)/2 y is number of assignments where x represents number of weeks in the course you want to find the number of weeks when the number of assignments is 30 (that is when y is 30) But it tells us to find the inverse function where the y and x are vise verse in meaning so you start by solving the following for y: x=(4y-3)/2 where y represents the number of weeks when x represents the number of assignments

OpenStudy (freckles):

So when you solve that for y. Plug in 30 for x to find y

OpenStudy (anonymous):

okay so correct me if I'm wrong because I'm pretty sure I am

OpenStudy (anonymous):

30 = (4y - 3)/2 and divide both sides by 2 15 = 4y - 3, then you add 3 from each side 18 = 4y, and divide by 4 y = 9/2

OpenStudy (freckles):

well that is right except I think they said they wanted you to find the inverse function

OpenStudy (freckles):

almost right *

OpenStudy (freckles):

2*30 is actually 60 not 15

OpenStudy (freckles):

so you would have 60=4y-3

OpenStudy (freckles):

but anyways i still think you are suppose to find the inverse function

OpenStudy (freckles):

30 = (4y - 3)/2 and *MULTIPLY* both sides by 2 but you want to solve x=(4y-3)/2 this first for y.

OpenStudy (anonymous):

wait so you multiply both sides by 2 and get 60 = 8y - 6/4? and then what?

OpenStudy (freckles):

?

OpenStudy (freckles):

no

OpenStudy (freckles):

I multiplied both sides by 2 60=4y-3

OpenStudy (freckles):

\[x=\frac{4y-3}{2} \\ 2 \cdot x =2 \cdot \frac{4y-3}{2} \\ \text{ you know that } \frac{2}{2}=1 \\ 2 \cdot x =\frac{2}{2} (4y-3) \\ 2x=1(4y-3) \\ 2x=4y-3\]

OpenStudy (anonymous):

and that's the inverse function?

OpenStudy (freckles):

no you have to solve for y

OpenStudy (freckles):

I did one step for you

OpenStudy (freckles):

I multiplied 2 on both sides

OpenStudy (freckles):

you did the other steps above the adding 3 on both sides then the dividing 4 on both sides do that here

OpenStudy (anonymous):

x = (4y - 3)/2 then you multiply both sides by 2 2x = 2/2(4y - 3) and since 2/2 = 1 we are left with 2x = 4y - 3 now all you have to do is plug 30 in for x to find y 2(30) = 4y - 3 60 = 4y - 3 add 3 to both sides 63 = 4y divide by 4 y = 15.75

OpenStudy (anonymous):

???

OpenStudy (freckles):

why don't you want to solve 2x=4y-3 for y?

OpenStudy (freckles):

2x=4y-3 add 3 to both sides 2x+3=4y then you divide 4 on both sides (2x+3)/4=y g(x)=(2x+3)/4 is the inverse function replace the number of assignments x with 30 to find the number of weeks y

OpenStudy (anonymous):

so (2(30)+3)/4?

OpenStudy (freckles):

yes you evaluate that which you already did without finding the inverse function --

OpenStudy (anonymous):

so the answer is y = 15.75?

OpenStudy (freckles):

well that is part of your solution

OpenStudy (freckles):

your problem asked to find the inverse function which I did for you above and then use your inverse function to find the number of weeks when the number of assignments is 30 and then part of your solution is also using complete sentences to describe how you found the inverse function

OpenStudy (anonymous):

okay, thank you